OpenStudy (anonymous):

the perimeter of a standard sized rectangular rug is 40ft. the length is 2ft longer than the width. find dimensions

OpenStudy (anonymous):

The equation of the perimeter of a rectangle is P = 2W + 2L. What is an equation relating the width, W, to the length, L?

OpenStudy (anonymous):

not understanding

OpenStudy (anonymous):

So what is the relationship between the length and the width?

OpenStudy (anonymous):

2ft longer than width not getting what your asking me

OpenStudy (anonymous):

Are you able to write this relationship as an equation?

OpenStudy (anonymous):

p=2*40+2*L

OpenStudy (anonymous):

is this what your asking

OpenStudy (anonymous):

The equation for your Perimeter is 40 = 2L + 2W. You know that the length is 2 ft longer than the width. The first step is to write an equation directly relating the length to the width.

OpenStudy (anonymous):

Hint: It involves the fact that the length is 2 ft greater than the width

OpenStudy (anonymous):

40=2*42+2*40

OpenStudy (anonymous):

No, don't use the equation for the perimeter. We need a second equation in order to solve that equation. Write a *new* equation, using the fact that the length is 2 ft longer than the width.

OpenStudy (anonymous):

p=2*42+2*40

OpenStudy (anonymous):

That is the equation for the perimeter. Since we know the length is 2 ft longer than the width, that means the L = W + 2. Now that we have a relationship between L and W, how can we use this new equation to solve for their dimensions?

OpenStudy (anonymous):

ok Im sorry that i dont get much of this im well into my 50s and im seriously trying

OpenStudy (anonymous):

L=40+2

OpenStudy (anonymous):

It's fine. Now that we have a relationship between the length and the width, we can substitute our new equation into the equation for our perimeter, P = 2L + 2W.

OpenStudy (anonymous):

P=2*2+2*40

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

Sorry, didn't get a notification. Since we know that L = W + 2, we can substitute that into our perimeter equation. This gives us 40 = 2* (W + 2) + 2W. Since we have this equation in terms of one variable, we can solve for that variable. This equation becomes 40 = 2W + 4 + 2W. Once we solve for W, we can solve for L, and thus solve this problem.